# Towards Automated Penetration Testing: Introducing LLM Benchmark, Analysis, and Improvements

### v3

@@ Abstract
Hacking poses a significant threat to cybersecurity, inflicting billions of dollars in damages annually. To mitigate these risks, ethical hacking, or penetration testing, is employed to identify vulnerabilities in systems and networks. Recent advancements in large language models (LLMs) have shown potential across various domains, including cybersecurity. However, there is currently no comprehensive, open, end-to-end automated penetration testing benchmark to drive progress and evaluate the capabilities of these models in security contexts. This paper introduces a novel open benchmark [^1] Benchmark has been shared: https://github.com/isamu-isozaki/AI-Pentest-Benchmark for LLM-based automated penetration testing, addressing this critical gap. We first evaluate the performance of LLMs, including GPT-4o and LLama 3.1-405B, using the state-of-the-art PentestGPT tool. Our findings reveal that while LLama 3.1 demonstrates an edge over GPT-4o, both models currently fall short of performing fully automated, end-to-end penetration testing. Next, we advance the state-of-the-art and present ablation studies that provide insights into improving the PentestGPT tool [^2] Our fork of PentestGPT with the modifications have been shared: https://github.com/isamu-isozaki/PentestGPT. Our research illuminates the challenges LLMs face in each aspect of Pentesting, e.g. enumeration, exploitation, and privilege escalation. This work contributes to the growing body of knowledge on AI-assisted cybersecurity and lays the foundation for future research in automated penetration testing using large language models.
@@ 1 Introduction
According to the 2023 Internet Crime Report by the Federal Bureau of Investigation (FBI), losses due to cybercrime totaled \$12.5 billion, reflecting a 20% increase from 2022 (Federal Bureau of Investigation, 2023). This upward trend highlights the escalating financial impact of cyber threats as digital reliance grows with the maturation of the internet age. Penetration testing, also referred to as ethical hacking or pen testing, is a critical security measure that involves simulating cyberattacks to identify system vulnerabilities (U.S. Department of the Interior, 2024). This approach helps organizations evaluate how well their systems can resist real-world attacks, uncovering potential weaknesses that attackers could exploit. While pen testing is essential for improving security and ensuring regulatory compliance, it cannot guarantee detection of all issues but effectively identifies the most common threats. Conducted by cybersecurity experts, these tests play a crucial role in mitigating risks and preventing costly breaches.
@@
We are in the midst of an AI revolution, with rapid advancements in Large Language Models (LLMs) opening up new possibilities across a wide spectrum of fields. In recent years, the field of AI has seen rapid advancements, with the seminal work of Vaswani (2017), which introduced transformers, driving much of the excitement around LLMs. The versatility and power of LLMs have prompted researchers and practitioners to explore their potential applications in nearly every domain of human knowledge and activity. Penetration testing is a task requiring deep expertise and extensive training which is currently being explored for potential automation through LLMs, which could significantly streamline the process Deng et al. (2024); Fang et al. (2024a, b, c); Happe et al. (2024). This shift towards AI-assisted penetration testing represents a paradigm change in how we approach cybersecurity assessments, potentially making them more accessible, efficient, and comprehensive. Our contributions in this paper are threefold. First, we introduce a novel benchmark to evaluate LLMs in the domain of penetration testing, filling a critical gap where no public benchmark previously existed. This benchmark aims to standardize the evaluation of AI models in cybersecurity contexts, facilitating more robust comparisons and driving progress in the field. Second, we assess this benchmark using the leading AI penetration testing tool, PentestGPT Deng et al. (2024), with two popular LLMs: GPT-4o and Llama3.1-405B Dubey et al. (2024). This assessment provides valuable insights into their performance, highlighting both the potential and current limitations of LLMs in cybersecurity applications. Third, we conduct ablation studies to analyze performance limitations and pinpoint areas where PentestGPT underperforms. Based on these findings, we propose adjustments to enhance the LLMs’ effectiveness in penetration testing tasks, paving the way for future improvements in AI-assisted cybersecurity.
@@ 3. Reads exploit and creates/runs a script using the exploit without prompting the agent
There has also been work to automate Privilege Escalation with no human intervention Happe et al. (2024). In their work, the explanation for how to perform each task was not given and there were only hints which were given as ablation. While LLMs were able to perform Privilege Escalation on the author’s benchmark well, the authors noticed the LLMs lack common sense reasoning, such as not utilizing passwords that were discovered or repeating the same commands Happe et al. (2024). Thus, we currently argue that for practical results in penetration testing with AI assistance, humans need to play a role.
Now, for to what extent humans should play a role, some research mentioned that full auto-penetration testing is not what Pentesters want due to the potential damage it can cause or potential exposure of attacks Happe and Cito (2023). In fact, the main part of Penetration testing that there is demand for automation for is information gathering/enumeration Happe and Cito (2023). However, this begs the question, are LLMs good at enumeration?
@@ 3 Benchmark
For this benchmark, we followed the method of PentestGPT Deng et al. (2024) as this was the only paper before us that attempted an end-to-end Penetration Testing Benchmark. However, we made 4 notable exceptions:
1. We used only Vulnhub boxes in the benchmark. Vulnhub provides free, downloadable virtual machines designed for penetration testing and security research, which makes it ideal for reproducible benchmarking (Vulnhub, 2012). In contrast, retired HackTheBox machines, used in PentestGPT paper, are paywalled, and some particular steps in the pen-testing process may require a VPN connection in certain regions, such as Europe based on our experience. Vulnhub’s free availability lowers the cost of benchmarking and enhances reproducibility. The Vulnhub boxes were sourced from a popular GitHub repository, CTF-Difficulty, which assigns difficulty ratings to each Vulnhub box Ignitetechnologies (2023). The initial walkthroughs were also gathered from this repository. Additionally, we included an easy box not listed in the repository, Funbox, which we classified as easy based on task numbers and their similarity to other easy boxes. All other walkthroughs not listed in the repository were found online and are referenced in the benchmark.
@@
3. Clear rules to minimize human involvement: In the PentestGPT benchmark, the extent of human involvement was not clearly defined. Our goal during the evaluation was to minimize human participation. However, certain steps, such as using BurpSuite PortSwigger (2024) and Wireshark Wireshark (2024), both GUI-based tools, required human interaction. Additionally, as we began evaluating PentestGPT, we found that the LLM’s instructions often assumed a human assistant to perform tasks, such as navigating websites to search for potential exploits, even when the HTML source code was available. To reduce human involvement, we established strict rules defining what actions humans were permitted to take. For example, PentestGPT did not make it clear when a task failure was determined by the authors. In our benchmark, to constrain the search space and maintain feasibility, we imposed a limit of five attempts per step. Moreover, PentestGPT did not specify what should be sent to the LLM when visiting websites. In our evaluation, we clearly state that the full HTML should be provided to the model.
4. Evaluate all tasks: As we wanted to be comprehensive, while Deng et al. (2024) stopped evaluating once a single task failed for every box, we evaluated all tasks. When a task failed, we provided the necessary commands along with the expected outcome, as outlined in our benchmark, ensuring consistency across trials. This approach allowed us to assess the performance of the LLMs across all task types. The full rules can be seen in the Appendix A.
@@
The task types and their categories were directly referenced from the PentestGPT paper by Deng et al. (2024). The extensive list of tasks and their categories can be referenced in Table 4 (See Appendix). While creating the benchmark, we manually assigned each task to corresponding task type based on the definition.
@@ 4.1 Experiment Setup
LLMs Used: As mentioned earlier, we evaluated our benchmarks on two popular LLMs: Llama3.1-405B from Meta (Dubey et al., 2024) and GPT-4o from OpenAI. Both models were tested using a 128K context length. A quantized Llama model using FP8 precision was selected for our study, ensuring consistency with the reference model.
@@ 4.3 Ablations
Figure 5: Three ablations were performed in this study: (a) Base PentestGPT (b) Ablation 1: Inject Summary - We maintain the summary and create a summary of past summaries to preserve the knowledge of progress made and maintain history. (c) Ablation 2: Structured Generation - Here we have updated the reasoning module to maintain a structured todo list instead of an unstructured Penetration Testing Tree (PTT). Ablation 2 includes the changes from Ablation 1. (d) Ablation 3: RAG Context - Building on Ablations 1 and 2, we add RAG context based on data scraped from Hacktricks Polop (2024). RAG retrieves similar chunks from the vectorDB to add to the context of the reasoning module.
@@ 4.3.2 Ablation 2: Structured Generation
For the PentestGPT method, the authors created a tree-like task structure called Penetration Testing Tree Deng et al. (2024). However, one issue with this approach is that this is only stored in natural language and there has been no processing to drop it down into a data structure like a list. We hypothesize that this leads to more hallucinations in the context of task planning. Thus, for this ablation, we moved to maintain a to-do list in the reasoning module (See Fig. 5c). To accomplish this, inspired by Wu et al. (2023), we used a ReAct agent Yao et al. (2022) tool calling-based approach to add useful tasks, remove unnecessary tasks, and finally modify the progress of each task to one of "done", "todo" or "in progress" where there can only be one in progress task. Exploring constrained generation techniques Willard and Louf (2023) for this stage was considered, but computational limitations in our current setup precluded its implementation within the timeframe of this research.
@@ 4.3.3 Ablation 3: Retrieval Augmented Generation
For this, we used Retrieval Augmented Generation Lewis et al. (2020) on the summary from the tool call stage for reference for adding new tasks (See Fig. 5d). We hypothesized this would be beneficial as in each pentest box, especially those in higher difficulty, there seems to be an increased focus on penetration testers going to the internet and researching exploits as opposed to using the knowledge they already have. For the database, we scraped the website contents from HackTricks Polop (2024) and chunked them into 500-word segments. These chunks were then indexed using the bge-large embedding model Xiao et al. (2023). The resulting embeddings are stored in ChromaDB Chroma (2024). During retrieval, we use cosine similarity between the summary and the chunks to select the top 3 documents. These top 3 are further refined to the top 2 using the bge-reranker Xiao et al. (2023).
@@ 5 Discussion
On the other hand, GPT 4o, even after 5 tries and we mention a method is not working, had the tendency to stick with a single task/rabbit hole and not give up on it where on occasion it kept repeating the same task over again which was also observed in other papers Happe et al. (2024).
@@ 6 Conclusion and Future work
We have found that at least for current LLM agents, even with human assistance for navigating websites/interpreting LLM commands, without help, were not able to complete a single end-to-end penetration testing experiment. Our analysis revealed that the two main categories where LLMs struggle are Reconnaissance, where LLama showed weakness, and Exploitation, which proved challenging for GPT-4. One area we are interested in pursuing is to increase the capability of our LLMs for Penetration Testing through Reinforcement Learning. We want to begin with a Penetration Testing Game with easier boxes, such as the boxes used in Happe, et al Happe et al. (2024). Another avenue we were interested in was to attempt to do self-play with LLMs to mirror human cybersecurity competitions, such as CCDC Competition (2024), where one agent attacks the network(red team) and the other defends(blue team) to progressively increase their capability.
@@ 8 Limitations
5. Ablation was done in full precision llama 3.1 405b which may give slightly different results than 8-bit precision which was used. However, studies have shown that 8-bit quantized LLama models have comparable performance to the full-weight counterparts (Li et al., 2024).
6. We only did one trial for each test, so the results may be more stochastic than doing multiple trials like in PentestGPT. However, we argue this was a trade-off as we went through every step in the benchmark instead of stopping once the test couldn’t progress, which was the case for PentestGPTDeng et al. (2024).
@@ Appendix D Categories and Task Types
The categories and tasks types used in this paper has been referenced from Deng et al. (2024). See Table 4.
@@
Table 4: Penetration Testing Categories and Task Types from Deng et al. (2024)
@@
Figure 12: Excerpt of prompts used. The prompts have been extended from Deng et al. (2024).

### v4

@@ Abstract
Hacking poses a significant threat to cybersecurity, inflicting billions of dollars in damages annually. To mitigate these risks, ethical hacking, or penetration testing, is employed to identify vulnerabilities in systems and networks. Recent advancements in large language models (LLMs) have shown potential across various domains, including cybersecurity. However, there is currently no comprehensive, open, automated, end-to-end penetration testing benchmark to drive progress and evaluate the capabilities of these models in security contexts. This paper introduces a novel open benchmark [^1] Benchmark has been shared: https://github.com/isamu-isozaki/AI-Pentest-Benchmark for LLM-based automated penetration testing, addressing this critical gap. We first evaluate the performance of LLMs, including GPT-4o and LLama 3.1-405B, using the state-of-the-art PentestGPT tool. Our findings reveal that while LLama 3.1 demonstrates an edge over GPT-4o, both models currently fall short of performing end-to-end penetration testing even with some minimal human assistance. Next, we advance the state-of-the-art and present ablation studies that provide insights into improving the PentestGPT tool [^2] Our fork of PentestGPT with the modifications have been shared: https://github.com/isamu-isozaki/PentestGPT. Our research illuminates the challenges LLMs face in each aspect of Pentesting, e.g. enumeration, exploitation, and privilege escalation. This work contributes to the growing body of knowledge on AI-assisted cybersecurity and lays the foundation for future research in automated penetration testing using large language models.
@@ 1 Introduction
According to the 2023 Internet Crime Report by the Federal Bureau of Investigation (FBI), losses due to cybercrime totaled \$12.5 billion, reflecting a 20% increase from 2022 [Federal Bureau of Investigation, 2023]. This upward trend highlights the escalating financial impact of cyber threats as digital reliance grows with the maturation of the internet age. Penetration testing, also referred to as ethical hacking or pen testing, is a critical security measure that involves simulating cyberattacks to identify system vulnerabilities [U.S. Department of the Interior, 2024]. This approach helps organizations evaluate how well their systems can resist real-world attacks, uncovering potential weaknesses that attackers could exploit. While pen testing is essential for improving security and ensuring regulatory compliance, it cannot guarantee detection of all issues but effectively identifies the most common threats. Conducted by cybersecurity experts, these tests play a crucial role in mitigating risks and preventing costly breaches.
@@
We are in the midst of an AI revolution, with rapid advancements in Large Language Models (LLMs) opening up new possibilities across a wide spectrum of fields. In recent years, the field of AI has seen rapid advancements, with the seminal work of Vaswani [2017], which introduced transformers, driving much of the excitement around LLMs. The versatility and power of LLMs have prompted researchers and practitioners to explore their potential applications in nearly every domain of human knowledge and activity. Penetration testing is a task requiring deep expertise and extensive training which is currently being explored for potential automation through LLMs, which could significantly streamline the process Deng et al. [2024], Fang et al. [2024a, b, c], Happe et al. [2024]. This shift towards AI-assisted penetration testing represents a paradigm change in how we approach cybersecurity assessments, potentially making them more accessible, efficient, and comprehensive. Our contributions in this paper are threefold. First, we introduce a novel benchmark to evaluate LLMs in the domain of penetration testing, filling a critical gap where no public benchmark previously existed. This benchmark aims to standardize the evaluation of AI models in cybersecurity contexts, facilitating more robust comparisons and driving progress in the field. Second, we assess this benchmark using the leading AI penetration testing tool, PentestGPT Deng et al. [2024], with two popular LLMs: GPT-4o and Llama3.1-405B Dubey et al. [2024]. This assessment provides valuable insights into their performance, highlighting both the potential and current limitations of LLMs in cybersecurity applications. Third, we conduct ablation studies to analyze performance limitations and pinpoint areas where PentestGPT underperforms. Based on these findings, we propose adjustments to enhance the LLMs’ effectiveness in penetration testing tasks, paving the way for future improvements in AI-assisted cybersecurity.
@@ 3. Reads exploit and creates/runs a script using the exploit without prompting the agent
There has also been work to automate Privilege Escalation with no human intervention Happe et al. [2024]. In their work, the explanation for how to perform each task was not given and there were only hints which were given as ablation. While LLMs were able to perform Privilege Escalation on the author’s benchmark well, the authors noticed the LLMs lack common sense reasoning, such as not utilizing passwords that were discovered or repeating the same commands Happe et al. [2024]. Thus, we currently argue that for practical results in penetration testing with AI assistance, humans need to play a role.
Now, for to what extent humans should play a role, some research mentioned that full auto-penetration testing is not what Pentesters want due to the potential damage it can cause or potential exposure of attacks Happe and Cito [2023]. In fact, the main part of Penetration testing that there is demand for automation for is information gathering/enumeration Happe and Cito [2023]. However, this begs the question, are LLMs good at enumeration?
@@ 3 Benchmark
For this benchmark, we followed the method of PentestGPT Deng et al. [2024] as this was the only paper before us that attempted an end-to-end Penetration Testing Benchmark. However, we made 4 notable exceptions:
1. We used only Vulnhub boxes in the benchmark. Vulnhub provides free, downloadable virtual machines designed for penetration testing and security research, which makes it ideal for reproducible benchmarking [Vulnhub, 2012]. In contrast, retired HackTheBox machines, used in PentestGPT paper, are paywalled, and some particular steps in the pen-testing process may require a VPN connection in certain regions, such as Europe based on our experience. Vulnhub’s free availability lowers the cost of benchmarking and enhances reproducibility. The Vulnhub boxes were sourced from a popular GitHub repository, CTF-Difficulty, which assigns difficulty ratings to each Vulnhub box Ignitetechnologies [2023]. The initial walkthroughs were also gathered from this repository. Additionally, we included an easy box not listed in the repository, Funbox, which we classified as easy based on task numbers and their similarity to other easy boxes. All other walkthroughs not listed in the repository were found online and are referenced in the benchmark.
@@
3. Clear rules to minimize human involvement: In the PentestGPT benchmark, the extent of human involvement was not clearly defined. Our goal during the evaluation was to minimize human participation. However, certain steps, such as using BurpSuite PortSwigger [2024] and Wireshark Wireshark [2024], both GUI-based tools, required human interaction. Additionally, as we began evaluating PentestGPT, we found that the LLM’s instructions often assumed a human assistant to perform tasks, such as navigating websites to search for potential exploits, even when the HTML source code was available. To reduce human involvement, we established strict rules defining what actions humans were permitted to take. For example, PentestGPT did not make it clear when a task failure was determined by the authors. In our benchmark, to constrain the search space and maintain feasibility, we imposed a limit of five attempts per step. Moreover, PentestGPT did not specify what should be sent to the LLM when visiting websites. In our evaluation, we clearly state that the full HTML should be provided to the model.
4. Evaluate all tasks: As we wanted to be comprehensive, while Deng et al. [2024] stopped evaluating once a single task failed for every box, we evaluated all tasks. When a task failed, we provided the necessary commands along with the expected outcome, as outlined in our benchmark, ensuring consistency across trials. This approach allowed us to assess the performance of the LLMs across all task types. The full rules can be seen in the Appendix A.
@@
The task types and their categories were directly referenced from the PentestGPT paper by Deng et al. [2024]. The extensive list of tasks and their categories can be referenced in Table 4 (See Appendix). While creating the benchmark, we manually assigned each task to corresponding task type based on the definition.
@@ 4.1 Experiment Setup
LLMs Used: As mentioned earlier, we evaluated our benchmarks on two popular LLMs: Llama3.1-405B from Meta [Dubey et al., 2024] and GPT-4o from OpenAI. Both models were tested using a 128K context length. A quantized Llama model using FP8 precision was selected for our study, ensuring consistency with the reference model.
@@ 4.3 Ablations
Figure 5: Three ablations were performed in this study: (a) Base PentestGPT (b) Ablation 1: Inject Summary - We maintain the summary and create a summary of past summaries to preserve the knowledge of progress made and maintain history. (c) Ablation 2: Structured Generation - Here we have updated the reasoning module to maintain a structured todo list instead of an unstructured Penetration Testing Tree (PTT). Ablation 2 includes the changes from Ablation 1. (d) Ablation 3: RAG Context - Building on Ablations 1 and 2, we add RAG context based on data scraped from Hacktricks Polop [2024]. RAG retrieves similar chunks from the vectorDB to add to the context of the reasoning module.
@@ 4.3.2 Ablation 2: Structured Generation
For the PentestGPT method, the authors created a tree-like task structure called Penetration Testing Tree Deng et al. [2024]. However, one issue with this approach is that this is only stored in natural language and there has been no processing to drop it down into a data structure like a list. We hypothesize that this leads to more hallucinations in the context of task planning. Thus, for this ablation, we moved to maintain a to-do list in the reasoning module (See Fig. 5c). To accomplish this, inspired by Wu et al. [2023], we used a ReAct agent Yao et al. [2022] tool calling-based approach to add useful tasks, remove unnecessary tasks, and finally modify the progress of each task to one of "done", "todo" or "in progress" where there can only be one in progress task. Exploring constrained generation techniques Willard and Louf [2023] for this stage was considered, but computational limitations in our current setup precluded its implementation within the timeframe of this research.
@@ 4.3.3 Ablation 3: Retrieval Augmented Generation
For this, we used Retrieval Augmented Generation Lewis et al. [2020] on the summary from the tool call stage for reference for adding new tasks (See Fig. 5d). We hypothesized this would be beneficial as in each pentest box, especially those in higher difficulty, there seems to be an increased focus on penetration testers going to the internet and researching exploits as opposed to using the knowledge they already have. For the database, we scraped the website contents from HackTricks Polop [2024] and chunked them into 500-word segments. These chunks were then indexed using the bge-large embedding model Xiao et al. [2023]. The resulting embeddings are stored in ChromaDB Chroma [2024]. During retrieval, we use cosine similarity between the summary and the chunks to select the top 3 documents. These top 3 are further refined to the top 2 using the bge-reranker Xiao et al. [2023].
@@ 5 Discussion
On the other hand, GPT 4o, even after 5 tries and we mention a method is not working, had the tendency to stick with a single task/rabbit hole and not give up on it where on occasion it kept repeating the same task over again which was also observed in other papers Happe et al. [2024].
@@ 6 Conclusion and Future work
We have found that at least for current LLM agents, even with human assistance for navigating websites/interpreting LLM commands, without help, were not able to complete a single end-to-end penetration testing experiment. Our analysis revealed that the two main categories where LLMs struggle are Reconnaissance, where LLama showed weakness, and Exploitation, which proved challenging for GPT-4. One area we are interested in pursuing is to increase the capability of our LLMs for Penetration Testing through Reinforcement Learning. We want to begin with a Penetration Testing Game with easier boxes, such as the boxes used in Happe, et al Happe et al. [2024]. Another avenue we were interested in was to attempt to do self-play with LLMs to mirror human cybersecurity competitions, such as CCDC Competition [2024], where one agent attacks the network(red team) and the other defends(blue team) to progressively increase their capability.
@@ 8 Limitations
5. Ablation was done in full precision llama 3.1 405b which may give slightly different results than 8-bit precision which was used. However, studies have shown that 8-bit quantized LLama models have comparable performance to the full-weight counterparts [Li et al., 2024].
6. We only did one trial for each test, so the results may be more stochastic than doing multiple trials like in PentestGPT. However, we argue this was a trade-off as we went through every step in the benchmark instead of stopping once the test couldn’t progress, which was the case for PentestGPTDeng et al. [2024].
@@ Appendix D Categories and Task Types
The categories and tasks types used in this paper has been referenced from Deng et al. [2024]. See Table 4.
@@
Table 4: Penetration Testing Categories and Task Types from Deng et al. [2024]
@@
Figure 12: Excerpt of prompts used. The prompts have been extended from Deng et al. [2024].